# Framing Unionization on Facebook: Communication around Representation Elections in the United States

### v1

@@ B.1 Licenses
The RoBERTa base model used for building the multi-label communication frames classifier is licensed under a MIT license.
Appendix C Robustness Checks
C.1 Window Size for Rolling Aggregation
The usage of frames is robust to the choice of the window size for rolling aggregation: despite minor changes, differences between won and lost representation cases are stable both pre- (cf. Figure C1) and post-elections (cf. Figure C2). For all plots, we balanced the number of wins and losses per union.
((a)) $n_{days}=5$
((b)) $n_{days}=3$
Figure C1: Usage of frames leading to elections, for lost and won cases. Results are reported in terms of mean and 95% confidence interval. Values on the y-axis refer to detrended scores. The $\star$ sign refers to the significance of the test for difference in means between the losing and winning: $\star$ indicates significance at $\alpha=0.05$, $\star\star$ at $\alpha=0.01$, and $\star\star\star$ at $\alpha=0.001$.
((a)) $n_{days}=5$
((b)) $n_{days}=3$
Figure C2: Usage of frames post-elections, for lost and won cases. Results are reported in terms of mean and 95% confidence interval. Values on the y-axis refer to detrended scores. The $\star$ sign refers to the significance of the test for difference in means between the losing and winning: $\star$ indicates significance at $\alpha=0.05$, $\star\star$ at $\alpha=0.01$, and $\star\star\star$ at $\alpha=0.001$.
C.2 Imbalance for Number of Cases per Union
@@
To assess robustness to the choice of random seed for sampling when balancing the number of wins and losses per union, we repeated the sampling across 20 random seeds. For each seed and frame, we performed independent t-tests comparing the difference in means between lost and won cases. Table C1 summarizes the results, reporting the percentage of seeds (over the 20 in total) for which the difference is significant at standard thresholds, along with median $t$-statistics, $p$-values, and ranges. Differences in all frames remain consistently significant.
Statistic Diagnostic Prognostic Motivational Community Engagement
% $p<0.05$ 100% 100% 100% 100% 100%
% $p<0.01$ 100% 100% 100% 100% 100%
% $p<0.001$ 100% 100% 100% 100% 100%
Median $t$ -11.11 -15.28 -7.17 -5.49 5.42
Median $p$ $2.2\times 10^{-28}$ $1.7\times 10^{-51}$ $8.6\times 10^{-13}$ $4.2\times 10^{-8}$ $6.1\times 10^{-8}$
Range $t$ [-12.53, -10.29] [-16.52, -14.43] [-8.03, -6.85] [-6.38, -4.11] [4.51, 6.44]
Table C1: t-test results for the difference in means between lost and won cases across 20 random seeds for balancing wins and losses per union. Reported are the percentage of tests below standard significance thresholds, the median $t$-statistic and $p$-value, and the range of $t$-statistics.
Next, we aimed at assessing the robustness to differences in the number of representation cases per union. We limited the total cases per union to the 90th percentile of the distribution (considering winning and losing outcomes jointly), and repeated the analyses across 20 random seeds. We first balanced the number of wins and losses per union for each seed. Then, for each seed and frame, we performed independent t-tests comparing the difference in means between lost and won cases. Table C2 summarizes the results, reporting the percentage of seeds (over the 20 in total) for which the difference is significant at standard thresholds, along with median $t$-statistics, $p$-values, and ranges. Differences in the diagnostic, prognostic, and community frames remain consistently significant, while motivational and engagement frames are not robust to the down-sampling procedure.
Statistic Diagnostic Prognostic Motivational Community Engagement
% $p<0.05$ 100% 100% 85% 100% 95%
% $p<0.01$ 100% 100% 25% 100% 65%
% $p<0.001$ 100% 100% 0% 90% 10%
Median $t$ -6.16 -7.95 -2.34 -3.66 2.66
Median $p$ $8.1\times 10^{-10}$ $2.5\times 10^{-15}$ $1.9\times 10^{-2}$ $2.5\times 10^{-4}$ $7.9\times 10^{-3}$
Range $t$ [-6.79, -5.44] [-8.46, -7.31] [-2.93, -1.42] [-4.31, -3.17] [1.85, 3.71]
Table C2: t-test results for the difference in means for lost and won cases across 20 random seeds. Reported are the percentage of tests below standard significance thresholds, the median $t$-statistic and $p$-value, and the range of $t$-statistics.
@@
Given that we balanced the number of wins and losses per union, we verified the robustness of results to the choice of random seed for sampling. As such, we repeated the balancing procedure with 20 random seeds. For each combination of pattern, frame and seed, we then computed the offset between lost and won cases (losing minus winning) and measured the significance of such an offset. In the end, we considered the average offset value across seeds for each pattern and frame combination and highlighted the offset as significant if 80% of the seeds showed a significant offset for that combination. Figure C3 summarizes these robustness results. The most relevant findings are robust: communication in won cases showed a higher frequency of decrease in prognostic and motivational frames after the election, and communication in lost cases was characterized by greater stability in the use of the motivational frame.
Figure C3: Usage of frames before and after the event, comparing lost and won cases. We report raw counts of cases assigned to each cell for losses (L) and wins (W). The $\star$ sign indicates a significant difference between outcome types at $\alpha=0.05$.
To account for potential imbalance in the number of representation cases per union, we limited the total cases per union to the 90th percentile, repeating the analyses across 20 random seeds. For each seed and frame, post-/pre-election changes were classified into three patterns: decrease, stable, or increase. For each seed, we first balanced the number of wins and losses per union. Then, for each combination of pattern, frame and seed, we computed the offset between lost and won cases (losing minus winning) and measured the significance of such an offset. In the end, we considered the average offset value across seeds for each pattern and frame combination and highlighted the offset as significant if 80% of the seeds showed a significant offset for that combination. Figure C4 summarizes these robustness results. While not all patterns remain significant under this procedure, the most relevant findings are robust: communication in won cases showed a higher frequency of decrease in prognostic and motivational frames after the election.
Figure C4: Usage of frames before and after the event, comparing lost and won cases. We report raw counts of cases assigned to each cell for losses (L) and wins (W). The $\star$ sign indicates a significant difference between outcome types at $\alpha=0.05$.
Appendix D Details on Labor Unions
Table D1 reports the acronyms and full names for national-level unions considered in the study. Figure D1 shows the distribution of the number of Facebook posts per labor union in the 2015-2024 period, highlighting some skewness.
@@ Appendix D Details on Labor Unions
Table D1: Union acronyms, expanded names, and organizational structure (Craft vs. Industrial).
Figure D1: Number of Facebook posts by labor union.

### v2

@@ Appendix B Modeling and Computational Details
We also evaluated a BERT-base uncased model that underwent the same fine-tuning pipeline and resulted in similar computational costs.
@@ B.1 Licenses
The RoBERTa base model used for building the multi-label communication frames classifier is licensed under a MIT license. The BERT-base model used as a comparison is licensed under an Apache License, Version 2.0. The emotion classification model is licensed under an Apache License, Version 2.0.
Appendix C Topic Modeling on Selected Unions
To explore the relationship between topics and frames, we examined posts from two of the most active (i.e., by total number of posts) craft and industrial unions on Facebook, namely IAFF and SEIU. Using topic modeling through BERTopic, we analyzed their posts in terms of prevalent topics. Figure C1 presents manually merged, overarching topics that appear in at least 5% of posts, showing how the prevalence of each topic within a frame deviates from its overall usage.
((a)) IAFF (Craft)
((b)) SEIU (Industrial)
Figure C1: Usage of topics within posts with a given frame relative to overall topic usage. Cell values represent percentages; red indicates higher-than-baseline usage, blue indicates lower-than-baseline usage.
In the main text, we pointed at the fact that applying topic modeling across all posts resulted in many union-specific topics. Even at the union level, some topics remain highly specific—for example, sharing safety tips for IAFF. Certain topics align with particular frames: in IAFF, union stories or events is more common in posts using the community frame, while sharing safety tips appears more often in the engagement frame. In SEIU, reasons to join a union and support for strikes are more often associated with community frames. At the same time, many topics, such as IAFF’s general info on union or SEIU’s rights for workers and fight for bill, cut across frames without a single dominant association.
Overall, topic modeling is useful for capturing additional nuance within frames, highlighting the specific issues discussed. However, its highly union-specific nature limits its usefulness for generalizable predictions beyond frame-level analysis.
Appendix D Robustness Checks
D.1 Alternative Classifiers
We considered two alternative classifiers in to compare with the RoBERTa-base one fine-tuned on language and then on labeled data, referenced in the main text. The first is a RoBERTa-base classifier simply fine-tuned on the labeled data, skipping the language fine-tuning step. We report its performance on the test set in Table D1.
Frame Precision Recall F1-score Support
Diagnostic 0.77 0.86 0.81 84
Prognostic 0.70 0.72 0.71 58
Motivational 0.59 0.83 0.69 29
Community 0.67 0.72 0.69 81
Engagement 0.72 0.82 0.77 51
Micro avg 0.70 0.79 0.74 303
Macro avg 0.69 0.79 0.73 303
Weighted avg 0.70 0.79 0.74 303
Table D1: Performance of the multi-label discourse frames classifier (RoBERTa-base, just fine-tuned for the classification task).
The second is a BERT-base uncased model, fine-tuned on language across all Facebook posts and then on the labeled data for multi-label classification. We report its performance in Table D2.
Frame Precision Recall F1-score Support
Diagnostic 0.80 0.82 0.81 84
Prognostic 0.67 0.62 0.64 58
Motivational 0.58 0.66 0.61 29
Community 0.62 0.65 0.63 81
Engagement 0.62 0.67 0.64 51
Micro avg 0.67 0.70 0.68 303
Macro avg 0.66 0.68 0.67 303
Weighted avg 0.67 0.70 0.68 303
Table D2: Performance of the multi-label discourse frames classifier (BERT-base uncased, fine-tuned for language and then fine-tuned for the classification task).
D.2 Coding Diagnostic and Prognostic in a Single Category
The predictive power of frames over election winning is robust to the coding of diagnostic and prognostic into a single category rather than considering them as separate predictors. The only exception is disgust, whose posterior distribution excludes zero in fewer than six out of ten model runs when diagnostic and prognostic frames are modeled separately, indicating sensitivity to model specification rather than a robust association. Figures D1 and D2 illustrate this pattern for models including only the proportion of diagnostic and prognostic posts prior to the election, respectively.
Figure D1: Posterior mean log-odds estimates across ten seeds, not considering the proportion of prognostic frames in the predictors. Variables not shown had no credible association with the outcome or credible association for fewer than six seeds.
Figure D2: Posterior mean log-odds estimates across ten seeds, not considering the proportion of diagnostic frames in the predictors. Variables not shown had no credible association with the outcome or credible association for fewer than six seeds.
D.3 Window Size for Aggregation
The estimated effects of diagnostic and prognostic frames on election outcomes remain stable across alternative aggregation window sizes. In contrast, the posterior distribution for community framing increasingly overlaps zero, indicating reduced evidence for an association under this specification. The estimated effect of neutral emotional content remains robust across window sizes. Random intercepts for states are largely stable, with the exception of Arizona, whose posterior distribution is shifted toward higher baseline odds of winning in this configuration (cf. Figure D3).
Figure D3: Posterior mean log-odds estimates across ten seeds, considering a window of three days for aggregation. Variables not shown had no credible association with the outcome or credible association for fewer than six seeds.
D.4 Imbalance for Number of Cases per Union
@@
In the main paper, we already demonstrated the robustness of coefficient estimates to different random seeds used to balance wins and losses within unions.
We next assess robustness to variation in the number of representation cases per union. Specifically, we capped the total number of cases per union at the 90th percentile of the distribution (considering winning and losing outcomes jointly) and repeated the analyses across ten random seeds. For each seed, we first balanced the number of wins and losses per union and then fitted a Bayesian logistic regression model predicting the binary election outcome.
Figure D4 presents posterior log-odds estimates for selected model variables. Overall, the results are robust to differences in the number of representation cases across unions. Estimated effects of diagnostic/prognostic and community frames remain stable, whereas posterior distributions for disgust-related and emotionally neutral content increasingly overlap zero under this specification. State-level random intercepts are largely consistent across models, with Colorado, Ohio, and Washington exhibiting similar deviations from the global intercept, while California’s posterior distribution is shifted toward higher baseline odds of winning.
Figure D4: Posterior mean log-odds estimates across ten seeds, downsampling over-represented unions.
@@
Given that we balanced the number of wins and losses per union, we verified the robustness of results to the choice of random seed for sampling. As such, we repeated the balancing procedure with 20 random seeds. For each combination of pattern, frame, and seed, we then computed the offset between lost and won cases (losing minus winning) and measured the significance of such an offset. In the end, we considered the average offset value across seeds for each pattern and frame combination and highlighted the offset as significant if 80% of the seeds showed a significant offset for that combination. Figure D5 summarizes these robustness results. The most relevant findings are robust: communication in won cases showed a higher frequency of decrease in prognostic and motivational frames after the election, and communication in lost cases was characterized by greater stability in the use of the motivational frame.
Figure D5: Usage of frames before and after the event, comparing lost and won cases. The $\star$ sign indicates a significant difference between outcome types ($p<0.05$).
Figure D6: Usage of frames before and after the event, comparing lost and won cases. The $\star$ sign indicates a significant difference between outcome types ($p<0.05$).
To account for potential imbalance in the number of representation cases per union, we limited the total cases per union to the 90th percentile, repeating the analyses across 20 random seeds. For each seed and frame, post-/pre-election changes were classified into three patterns: decrease, stable, or increase. For each seed, we first balanced the number of wins and losses per union. Then, for each combination of pattern, frame and seed, we computed the offset between lost and won cases (losing minus winning) and measured the significance of such an offset. In the end, we considered the average offset value across seeds for each pattern and frame combination and highlighted the offset as significant if 80% of the seeds showed a significant offset for that combination. Figure D6 summarizes these robustness results. While not all patterns remain significant under this procedure, the most relevant findings are robust: communication in won cases showed a higher frequency of decrease in prognostic and motivational frames after the election.
Appendix E Details on Labor Unions
Table E1 reports the full names unions considered in the study. Figure E1 shows the distribution of the number of Facebook posts per union.
@@ Appendix E Details on Labor Unions
Table E1: Union acronyms, expanded names, and organizational structure (Craft vs. Industrial).
Figure E1: Number of Facebook posts by labor union.